ap stats probability

0.0(0)
Studied by 0 people
call kaiCall Kai
learnLearn
examPractice Test
spaced repetitionSpaced Repetition
heart puzzleMatch
flashcardsFlashcards
GameKnowt Play
Card Sorting

1/17

encourage image

There's no tags or description

Looks like no tags are added yet.

18 Terms

1
New cards

sample space

the complete set of all possible, non-overlapping outcomes of a random process or chance experiment.

  • roll a die S={1,2,3,4,5,6}


2
New cards

event

any collection of outcomes from the sample space

  • rolling a prime # E={2,3,5}

  • rolling a prime # or even number E = {2,3,4,5,6}


3
New cards

simple event

an event consisting of exactly one outcome

  • rolling a “2” E = {2}


4
New cards

complement

for any event E, P(E) + P(not E) = 1

5
New cards

independent events

two events are independent if knowing that one will occur/has occurred does not change the probability that the other occurs.

+ must be able to occur simultaneously, cannot be mutually exclusive

  1. a randomly selected student is a senior - what is the probability they play soccer for PWSH?

  • likely independent

  1. a randomly selected student is a junior - what is the probability they will graduate this school year?

  • dependent


6
New cards

checking for independence

  1. P(A and B) = P(A) * P(B)

  2. P(A|B) = P(A)

  3. P(B|A) = P(B)


+ P(A|B) = P(A) = P(A|Bc)


7
New cards

mutually exclusive/disjoint

if events are mutually exclusive, only one can occur at a time

+ two events have no outcomes in common (ex. rolling a 2 and a 5 on the same roll)

P(A and B) = 0


8
New cards

probability multiplication rules

if two events are indep: P(A and B) = P(A) * P(B)

not indep, use general rule: P(A and B) = P(A) * P(B|A)

9
New cards

probability addition rules

if two events are mutually exclusive: P(E or F) = P(E) + P(F)

if not, use general addition rule: P(E or F) = P(E) + P(F) - P(E and F)

indep events version. : P(E or F) = P(E) + P(F) - [P(E) * P(F)]

10
New cards

exactly one example

light bulb defective 5% of time. you randomly pick a package of two such bulbs off the shelf of a store. what is the probability that exactly one bulb is defective?


P(exactly one) = P(D and Dc) or P(Dc and D)

= (0.95)(0.95) + (0.95)(0.05)

= (0.095)

11
New cards

at least one

P(at least one) = 1 - P(Dc and Dc)


same example above, but probability that at least one bulb is defective

P(at least one) = 1 - P(0.95)²

= 0.0975.

12
New cards

conditional probability

a probability that takes into account a given condition

P(B|A) = P(A and B) / P(A)

or = P(and)/P(given)

13
New cards

law of large numbers

as the number of repetitions of a chance experiment increase, the difference between the relative frequency of occurrence for an event and the true probability approaches zero

  • with larger n, experimental probability approaches the theoretical probability


14
New cards

simulation summary and drawing conclusions only appropriate

~at least 100 trials

15
New cards

iqr + outlier formulas

  • Interquartile Range: IQR = Q₃ - Q₁

  • Lower Outlier Fence: Lower Boundary = Q₁ - (1.5 × IQR)

  • Upper Outlier Fence: Upper Boundary = Q₃ + (1.5 × IQR)



SD, mean-based

lower: mean - 2s

uppwer: mean + 2s

16
New cards

sample variance formula

knowt flashcard image
17
New cards

population variance formula

knowt flashcard image
18
New cards